OpenStudy (anonymous):

Line AB contains points A (1, 2) and B (−2, 6). The slope of line AB is

OpenStudy (cpinder):

y2 - y1 over x2 - x1 will give you your slope.

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

so 6-2=4 -2-1=-3 which gives me 3/4

OpenStudy (anonymous):

whixh makes it a positive slope right?

OpenStudy (anonymous):

wait no 4/-3 so a negative slope

OpenStudy (anonymous):

yea your correct.. good job @ashhbashh17